You throw a ball vertically upward from the roof of a tall building. The ball leaves your hand at a

point even with the roof railing with an upward speed of 15.0 m/s; the ball is then in free fall. On its

way back down, it just misses the railing. Find (a) the ball’s position and velocity 1.00 s and 4.00 s

after leaving your hand; (b) the ball’s velocity when it is 5.00 m above the railing; (c) the maximum

height reached; (d) the ball’s acceleration when it is at its maximum height.


1
Expert's answer
2021-10-31T18:16:00-0400

a)

"h=vt-\frac{gt^2}2,"

"h_1=15-5=10~m,"

"h_4=60-80=-20~m,"

"v=v_0-gt,"

"v_1=15-10=5~\frac ms,"

"v_4=15-40=-25~\frac ms,"

b)

"v=±\sqrt{2gh}=±10~\frac ms,"

c)

"h_{max}=\frac{v^2}{2g}=11.3~m,"

d)

"a(h_{max})=g=10~\frac m{s^2}."